# The formation of Schrödinger cat-like states in the process of spontaneous parametric down-conversion

### v5

@@ IV.3 Robustness of super-Poissonian statistics to dissipation
The Fano factor $FF_{j}$ characterizes photon statistics: $FF_{j}<1$ indicates sub-Poissonian, $FF_{j}=1$ Poissonian, and $FF_{j}>1$ super-Poissonian behavior. At $\tau=0.38$, the calculated values from Eq. (8) are: $FF_{1}=8.54$ (non-diss.), $8.41$ (diss.); $FF_{2}=6.67$ (non-diss.), $6.48$ (diss.)—all indicating super-Poissonian statistics.
IV.4 Entanglement of SCLSs
To quantify the entanglement between the two modes $(\hat{a}_{1},\hat{a}_{2})$ in the SCLSs, we compute the Schmidt number [Law2004] defined as:
@@ IV.4 Entanglement of SCLSs
where $\hat{\rho}_{1}$ and $\hat{\rho}_{2}$ are the reduced density matrices of modes $\hat{a}_{1}$ and $\hat{a}_{2}$, respectively. For pure bipartite states, the parameter $K$ serves as an entanglement measure, with $K>1$ indicating non-separable (entangled) states. At $\tau=0.38$, the value calculated from Eq. (9) is $K=1.93$ for the non-dissipative case, confirming the presence of entanglement.
V Summary and conclusions
@@
We demonstrate the formation of entangled Schrödinger cat-like states (SCLSs) in both fundamental and second-harmonic modes during spontaneous parametric down-conversion (SPDC) with quantum-depleted pumps, under both non-dissipative ($\gamma_{1,2}=0$) and dissipative ($\gamma_{1,2}=0.15$) regimes. At $\tau=0.38$, Wigner functions reveal clear SCLS formation with mode $\hat{a}_{1}$ exhibiting $N_{1}\approx 25.2$ and mode $\hat{a}_{2}$ showing $N_{2}\approx 7.3$ in the non-dissipative case, while maintaining non-Gaussianity (Wigner negativity) under dissipation. The dissipative regime introduces odd photon numbers into the SCLSs (see Figs. 3(c)-3(d)).
@@
These states demonstrate significant squeezing in quadrature components: 3.37-3.37 dB ($\Delta^{2}p_{1}=0.23$) for mode $\hat{a}_{1}$ and 3.76-3.76 dB ($\Delta^{2}x_{2}=0.21$) for mode $\hat{a}_{2}$, along with super-Poissonian statistics confirmed by Fano factor analysis. The super-Poissonian character ($FF\gg 1$) enhances correlation-based imaging through photon bunching [a2017], making these states valuable for quantum sensing [Salykina2023, Singh2024] applications, particularly in Mach-Zehnder interferometry [Shukla2023, Toscano2006, Salykina2023, Singh2024] and optical qubit encoding [Yamamoto2020, Lloyd1999].
@@
Odd SCLSs can be generated by changing the initial state to $|\psi_{0}\rangle=|1\rangle_{1}\otimes|\alpha_{20}\rangle_{2}$, producing only odd photon numbers in mode $\hat{a}_{1}$: $e^{-itg\hat{H}_{int}}|1\rangle_{1}\otimes|\alpha_{20}\rangle_{2}\approx|1\rangle_{1}\otimes|\alpha_{20}\rangle_{2}-i2tg\alpha_{20}\sqrt{6}|3\rangle_{1}\otimes|\alpha_{20}\rangle_{2}$. The analysis for odd SCLSs follows the same methodology as for even states and is omitted here. Recent attempts to find optimal generation timing for SCLSs remain an active research focus [Gorshenin2024].
@@
A key finding is the formation of entanglement between the modes, a direct consequence of the fully quantized treatment of the pump. This entanglement, combined with the non-Gaussian characteristics of the SCLSs, creates a powerful resource for quantum technologies. Furthermore, the relationship between the observed non-classicality and Bell-type non-classicality, which generally differ in nature [Volovich2016], presents an interesting avenue for further investigation.
@@
These states are particularly promising for enhancing the sensitivity of Mach-Zehnder interferometers [Shukla2023, Toscano2006, Salykina2023, Singh2024] and for applications in correlation-based imaging. From a practical perspective, these entangled non-Gaussian states can be generated in standard SPDC setups using periodically poled nonlinear crystals [see Belinsky2018, and references therein], requiring approximately 64% pump conversion efficiency. Their verification via homodyne tomography [Paul2024] is feasible and offers an advantage over $\chi^{(3)}$-based schemes due to lower pump intensities.
@@
Our findings not only contribute to the toolbox of continuous-variable quantum information processing [Holevo2019] but also open new avenues for exploring the interplay between entanglement, non-Gaussianity, and non-classicality in fundamental quantum optical processes.
Acknowledgements.
We thank Prof. Anatoly V. Masalov for discussions on obtained results and providing valuable inputs.

### v6

@@ 4.3 Robustness of super-Poissonian statistics to dissipation
The Fano factor $FF_{j}$ characterizes photon statistics: $FF_{j}<1$ indicates sub-Poissonian, $FF_{j}=1$ Poissonian, and $FF_{j}>1$ super-Poissonian behavior. At $\tau=0.38$, the calculated values from Eq. (8) are: $FF_{1}\approx 8.54$ (non-diss.), $8.35$ (diss.); $FF_{2}\approx 6.66$ (non-diss.), $6.38$ (diss.)—all indicating super-Poissonian statistics.
4.4 Entanglement of SCLSs
To quantify the entanglement between the two modes $(\hat{a}_{1},\hat{a}_{2})$ in the SCLSs, we compute the Schmidt number [34] defined as:
@@ 4.4 Entanglement of SCLSs
where $\hat{\rho}_{1}$ and $\hat{\rho}_{2}$ are the reduced density matrices of modes $\hat{a}_{1}$ and $\hat{a}_{2}$, respectively. For pure bipartite states, the parameter $K$ serves as an entanglement measure, with $K>1$ indicating non-separable (entangled) states. At $\tau=0.38$, the value calculated from Eq. (9) is $K\approx 1.93$ for the non-dissipative case, confirming the presence of entanglement.
5 Non-degenerate SPDC: An Extended Analysis
While the degenerate SPDC process is foundational, the non-degenerate (NDSPDC) configuration—where a pump photon splits into two distinguishable photons—is of paramount importance for quantum information processing, enabling heralded single-photon generation and discrete-variable entanglement. In this section, we adapt our fully quantum numerical methodology to the NDSPDC process to investigate the formation and properties of non-Gaussian states and entanglement in contrast to the degenerate case.
5.1 Hamiltonian and System Description
The interaction Hamiltonian for the NDSPDC process, in which a pump photon splits into two distinguishable signal and idler photons ($\omega_{p}=\omega_{s}+\omega_{i}$) propagating collinearly in a $\chi^{(2)}$ nonlinear crystal, is given by  [25, 26, 3, 27, 28]
$\hat{H}_{\mathrm{int}}^{(\mathrm{ND})}=\hbar g^{(\mathrm{ND})}(\hat{a}_{s}\hat{a}_{i}\hat{a}_{p}^{\dagger}+\mathrm{h.c.}),$ (10)
where $\hat{a}_{p}$, $\hat{a}_{s}$, and $\hat{a}_{i}$ are the annihilation operators for the pump, signal, and idler modes at frequencies $\omega_{p}$, $\omega_{s}$, and $\omega_{i}$, respectively. We set $\hbar=1$ for simplicity. The nonlinear coupling constant $g^{(\mathrm{ND})}$ differs from its degenerate counterpart $g^{(\mathrm{D})}$ due to distinct phase-matching conditions inherent to the two processes. The open-system dynamics are governed by a Lindblad master equation of the same form as Eq. (2), with $\hat{H}_{\mathrm{int}}$ replaced by $\hat{H}_{\mathrm{int}}^{(\mathrm{ND})}$ and with dissipation channels for all three modes.
For this brief comparative study, we restrict the NDSPDC analysis to the non-dissipative regime ($\gamma_{j}=0$) to examine the ideal, lossless quantum dynamics. We choose an initial state analogous to the degenerate case: the signal and idler modes in the vacuum state and the pump mode in a coherent state,
$|\psi_{0}^{(\mathrm{ND})}\rangle=|0\rangle_{s}\otimes|0\rangle_{i}\otimes|\alpha_{p0}\rangle_{p},$ (11)
with $|\alpha_{p0}|^{2}=20$ and phase $\varphi_{p0}=\pi/2$.
5.2 Mean number of photons and formation of Non-Gaussian States
We compute the mean photon numbers for both the signal ($\hat{a}_{s}$) and pump ($\hat{a}_{p}$) modes using Eq. (4), and their corresponding Wigner functions $W_{s}^{(\mathrm{ND})}(\alpha_{s},\tau)$ and $W_{p}^{(\mathrm{ND})}(\alpha_{p},\tau)$ using Eq. (3) from the reduced density matrices $\hat{\rho}_{s}^{(\mathrm{ND})}(\tau)$ and $\hat{\rho}_{p}^{(\mathrm{ND})}(\tau)$. The resulting phase-space portraits, shown in Figs. 4(a) and 4(b), reveal the formation of non-Gaussian quantum states.
The pump mode’s Wigner function [Fig. 4(b)] exhibits clear interference fringes and negative-value regions, a definitive signature of Schrödinger cat-like state (SCLS) formation. This parallels the behavior observed for the pump mode in the degenerate SPDC process, confirming that a fully quantized pump leads to non-Gaussian state generation in both $\chi^{(2)}$ processes.
In contrast to the pump, the signal mode’s Wigner function [Fig. 4(a)] displays a markedly different structure. The semiclassical, undepleted-pump approximation for SPDC predicts the signal mode to evolve into a thermal (Gaussian) state. However, under the full quantum treatment with pump depletion, we observe the formation of a non-Gaussian state. This deviation thus demonstrates that a fully quantized pump model is required to produce non-Gaussian states in the down-converted fields. Due to the symmetry of the Hamiltonian in Eq. (10), the idler mode’s Wigner function $W_{i}^{(\mathrm{ND})}(\alpha_{i},\tau)$ is qualitatively identical to that of the signal mode; therefore, its explicit analysis is omitted.
Figure 4: Wigner functions (a) $W_{s}^{ND}(\alpha_{s},\tau)$ and (b) $W_{p}^{ND}(\alpha_{p},\tau)$ for non-degenerate SPDC process at $\tau=1$, shown for the non-dissipative case ($\gamma_{1,2,3}=0$). Mean photon numbers: $N_{s}\approx 5.44$, $N_{p}\approx 14.55$. Initial conditions: $\hat{a}_{s}$ and $\hat{a}_{i}$ in vacuum, $\hat{a}_{p}$ in a coherent state ($|\alpha_{p0}|^{2}=20$, $\varphi_{p0}=\pi/2$).
5.3 Entanglement Structure in NDSPDC
A fundamental distinction from the two-mode degenerate SPDC process is the tripartite nature of the state generated by NDSPDC. To quantify the bipartite entanglement within this system, we compute the Schmidt number for a specific partition. The entanglement between the signal mode and the combined idler-pump subsystem is
$K^{(\mathrm{ND})}(\tau)=\frac{1}{\operatorname{Tr}[(\hat{\rho}_{s}^{(\mathrm{ND})}(\tau))^{2}]}=\frac{1}{\operatorname{Tr}[(\hat{\rho}_{ip}^{(\mathrm{ND})}(\tau))^{2}]},$ (12)
where $\hat{\rho}_{ip}^{(\mathrm{ND})}$ is the reduced density matrix for the idler and pump modes. For the specified initial state and parameters, we find $K^{(\mathrm{ND})}\approx 10.38$ at $\tau=1.0$ in the non-dissipative case, indicating strong entanglement. This value is notably larger than the Schmidt number $K\approx 1.93$ found for the bipartite entanglement in the degenerate SPDC process at its characteristic time ($\tau=0.38$), reflecting the higher-dimensional entanglement structure accessible in the non-degenerate system.
6 Numerical implementation and convergence
To ensure the reliability and consistency of the numerical simulations, particular care was taken in the truncation of the Hilbert space and in monitoring several internal consistency checks. Since the underlying Hilbert space is infinite dimensional, all numerical calculations were performed in a truncated Fock basis with a sufficiently large cutoff $N_{\mathrm{max}}$. The value of $N_{\mathrm{max}}$ was increased until all relevant physical observables converged and became insensitive to further increases of the cutoff.
As consistency checks, we systematically monitored the following quantities throughout the evolution: (i) the trace of the density matrix, which was verified to remain equal to unity within numerical precision; (ii) the purity in the case of unitary dynamics, quantified by $\mathrm{Tr}(\rho^{2})$, which equals one for the full (pure) density matrix and satisfies $\mathrm{Tr}(\rho^{2})\leq 1$ for reduced density matrices obtained via partial tracing; (iii) the positivity of the density matrix, verified by confirming the non-negativity of its eigenvalues; (iv) the mean photon numbers and their variances, which were checked for convergence with respect to the Hilbert-space cutoff; and (v) the stability of phase-space representations, in particular the Wigner functions, whose qualitative features and interference patterns were found to be unchanged upon increasing $N_{\mathrm{max}}$.
These combined checks ensure that the numerical results are well converged and that truncation effects do not affect the physical conclusions presented in this paper.
7 Summary and conclusions
@@
We demonstrate the formation of entangled Schrödinger cat-like states (SCLSs) in both fundamental and second-harmonic modes during degenerate spontaneous parametric down-conversion (SPDC) with quantum-depleted pumps, under both non-dissipative ($\gamma_{1,2}=0$) and dissipative ($\gamma_{1,2}=0.15$) regimes. At $\tau=0.38$, Wigner functions reveal clear SCLS formation with mode $\hat{a}_{1}$ exhibiting $N_{1}\approx 25.28$ and mode $\hat{a}_{2}$ showing $N_{2}\approx 7.35$ in the non-dissipative case, while maintaining non-Gaussianity (Wigner negativity) under dissipation. The dissipative regime introduces odd photon numbers into the SCLSs (see Figs. 3(c)-3(d)).
@@
These states demonstrate significant squeezing in quadrature components: $4.03$ dB ($\Delta^{2}p_{1}\approx 0.1976$) for mode $\hat{a}_{1}$ and $4.04$ dB ($\Delta^{2}x_{2}\approx 0.1970$) for mode $\hat{a}_{2}$, along with super-Poissonian statistics confirmed by Fano factor analysis. The super-Poissonian character ($FF\gg 1$) enhances correlation-based imaging through photon bunching [35], making these states valuable for quantum sensing [10, 11] applications, particularly in Mach-Zehnder interferometry [8, 9, 10, 11] and optical qubit encoding [13, 36].
@@
Odd SCLSs can be generated by changing the initial state to $|\psi_{0}\rangle=|1\rangle_{1}\otimes|\alpha_{20}\rangle_{2}$, producing only odd photon numbers in mode $\hat{a}_{1}$: $e^{-it\hat{H}_{int}}|1\rangle_{1}\otimes|\alpha_{20}\rangle_{2}\approx|1\rangle_{1}\otimes|\alpha_{20}\rangle_{2}-itg\alpha_{20}\sqrt{6}|3\rangle_{1}\otimes|\alpha_{20}\rangle_{2}$. The analysis for odd SCLSs follows the same methodology as for even states and is omitted here. Recent attempts to find optimal generation timing for SCLSs remain an active research focus [37].
Extending our analysis to the non-degenerate SPDC (NDSPDC) process, we find that a fully quantized pump similarly leads to non-Gaussian state formation. The pump mode develops SCLS features, while the signal mode exhibits a non-Gaussian structure distinct from the thermal state predicted by the semiclassical approximation. Notably, the tripartite NDSPDC system exhibits a higher-dimensional entanglement structure, with a Schmidt number of $K^{(\mathrm{ND})}\approx 10.38$, compared to the bipartite entanglement ($K\approx 1.93$) of the degenerate process.
@@
A key finding is the formation of entanglement between the modes, a direct consequence of the fully quantized treatment of the pump. This entanglement, combined with the non-Gaussian characteristics of the SCLSs, creates a powerful resource for quantum technologies. Furthermore, the relationship between the observed non-classicality and Bell-type non-classicality, which generally differ in nature [38], presents an interesting avenue for further investigation.
@@
These states are particularly promising for enhancing the sensitivity of Mach-Zehnder interferometers [8, 9, 10, 11] and for applications in correlation-based imaging. From a practical perspective, these entangled non-Gaussian states can be generated in standard SPDC setups using periodically poled nonlinear crystals [see 5, and references therein], which requires high pump conversion efficiencies (approximately 63.2% for the degenerate case and similarly high for NDSPDC). Their verification via homodyne tomography [39] is feasible and offers an advantage over $\chi^{(3)}$-based schemes due to lower pump intensities.
@@
Our findings not only contribute to the toolbox of continuous-variable quantum information processing [40] but also open new avenues for exploring the interplay between entanglement, non-Gaussianity, and non-classicality in fundamental quantum optical processes.
Acknowledgements
We thank Prof. Anatoly V. Masalov for discussions on the obtained results and for providing valuable inputs. We also thank the Referee and Editor, Prof. Matteo G. A. Paris, for the insightful suggestions, including to extend our fully quantized treatment to the non-degenerate down-conversion configuration. Implementing this extension has allowed us to demonstrate the generation of non-Gaussian states and complex entanglement in a tripartite system, thereby broadening the scope of our work.